OpenStudy (anonymous):

what is the focus of the parabola given by the equation y=x^2-2x-3? a. (4,-15/4) b. (1,4) c. (1,-4) d. (1,15/4) e. (1,-15/4)

OpenStudy (misty1212):

i swear i just saw this question a second ago

OpenStudy (misty1212):

trick is to write it as \[y+3=x^2-2x\] then complete the square on the left do you know how to do that?

OpenStudy (misty1212):

lol i mean on the RIGHT

OpenStudy (anonymous):

no i dont :( @misty1212

OpenStudy (misty1212):

ok ready? it is not hard

OpenStudy (anonymous):

ready @misty1212

OpenStudy (misty1212):

what is half of 2?

OpenStudy (misty1212):

...

OpenStudy (anonymous):

1 @misty1212

OpenStudy (misty1212):

ok and \(1^2=1\) so we go right from \[y+3=x^2-2x\] to \[y+3+1=(x-1)^2\] or \[y+4=(x-1)^2\]

OpenStudy (misty1212):

now we need the focus, but first we need the vertex the vertex we see from our eyeballs is \((1,-4)\)

OpenStudy (misty1212):

then we compare to the standard form \[4p(y-k)=(x-h)^2\] and see that in this case \(4p=1\)so \(p=\frac{1}{4}\)

OpenStudy (misty1212):

so we travel \(\frac{1}{4}\) units up from \((1,-4)\) and get to \((1,-\frac{15}{4})\)

OpenStudy (misty1212):

what to check it

OpenStudy (anonymous):

yes please thank you @misty1212